OpenStudy (anonymous):

solve for x:11^x=1331

jimthompson5910 (jim_thompson5910):

11^x = 1331 11^x = 11^3 Since the bases are equal (to 11), the exponents are equal. So x = 3

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

You're welcome